OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (anonymous):

i dont get it , which is the answer

OpenStudy (lieutenantgeneral):

Where is says: Solutions: and look at your answer choices and see

OpenStudy (lieutenantgeneral):

which is correct

OpenStudy (anonymous):

oh , well I dont have an answer choice its a free response thats why I was confused

Parth (parthkohli):

Is it\[7\sin^2 x - 14\sin x + 2 = -5\]

OpenStudy (anonymous):

@ilovereyvis_x3 is it sin (2x) or (sin x)^2